# CURE: A dataset for Clinical Understanding & Retrieval Evaluation

### v3

@@ Abstract
Given the dominance of dense retrievers that do not generalize well beyond their training dataset distributions, domain-specific test sets are essential in evaluating retrieval. There are few test datasets for retrieval systems intended for use by healthcare providers in a point-of-care setting. To fill this gap we have collaborated with medical professionals to create CURE, an ad-hoc retrieval test dataset for passage ranking with 2000 queries spanning 10 medical domains with a monolingual (English) and two cross-lingual (French/Spanish $\rightarrow$ English) conditions. In this paper, we describe how CURE was constructed and provide baseline results to showcase its effectiveness as an evaluation tool. CURE is published with a Creative Commons Attribution Non Commercial 4.0 license and can be accessed on Hugging Face.
CURE: A dataset for Clinical Understanding & Retrieval Evaluation
Nadia Sheikh 1, Anne-Laure Jousse1, Daniel Buades Marcos1, Akintunde Oladipo1, Olivier Rousseau1, Jimmy Lin2 1Clinia, 2University of Waterloo, {nadia.sheikh, anne-laure.jousse, daniel.buades, akintunde.oladipo, olivier.rousseau}@clinia.com jimmylin@uwaterloo.ca
1 Introduction
@@
Evidence-based medicine requires point-of-care healthcare practitioners to keep abreast of advancements in their specialties so they can make data-driven decisions when providing care to patients Sackett (1997). Despite this imperative, it has been estimated that research findings can take up to 17 years to be fully implemented into practice. Balas and Boren (2000).
@@
One impediment for practitioners is the increasingly rapid rate at which relevant literature is published Kamtchum-Tatuene and Zafack (2021). Each year, there are over 1 million new articles published in the field of biomedicine and life sciences González-Márquez et al. (2024).
@@
Given the ever expanding biomedical information landscape, systems that support ad hoc retrieval are an essential tool for health care practitioners to provide evidence-based care Lu (2011). Using ad hoc retrieval, a health-care professional can express their information need as a query, initiating a search that returns likely relevant documents from a vast corpus Lu (2011).
@@
Although the vast majority of published biomedical literature is in English, healthcare providers may practice in other languages Baethge (2008). To effectively accommodate this linguistic diversity, tools aimed at efficient biomedical information access should enable cross-lingual retrieval.
@@
Information retrieval systems and techniques are typically evaluated using test datasets. There are a wide variety of general test datasets available for evaluating retrieval. At its inception, MTEB (Massive Text Embedding Benchmark) made available 15 retrieval test sets, a collection that continues to expand Muennighoff et al. (2022). To the best of our knowledge, there are no cross-lingual datasets suitable for evaluating retrieval systems intended for use by healthcare providers in a point-of-care setting. As such, domain-specific test sets are essential given the dominance of dense retrievers, whose performance does not generalize to out-of-domain tasks Thakur et al. (2021); Ren et al. (2023).
@@
At Clinia, we build products that power search and discovery in health workflows, and the absence of such test datasets has hampered product development, as we cannot benchmark developed solutions systematically. We are working to close this gap by introducing a domain-specific test set which we have already released under an open-source license.
@@
In this work, we introduce Clinical Understanding & Retrieval Evaluation (CURE), a test dataset to evaluate ad hoc retrieval systems for use by medical practitioners in a point-of-care setting. We present results from competitive baselines to demonstrate the utility of CURE as a test dataset and enable comparison with future systems.
@@
CURE is a retrieval dataset with a monolingual and two cross-lingual conditions, with splits spanning ten medical domains. Queries in CURE are natural language questions formulated by health care providers. They express the information needs of practitioners consulting academic literature in the course of their duties. Queries are available in English, French and Spanish. The corpus is constructed by mining an index of english passages extracted from biomedical academic articles. We have made this dataset available in an effort to spur research and development on retrieval in a point-of-care setting.
2 Dataset Construction
@@
CURE is a test collection for passage ranking. In passage ranking, given a corpus C of n passages and a query q, the objective is to retrieve a list of k passages, sorted by decreasing relevance. A relevant passage is one that meets the information need expressed by the query. In the context of a dataset, the relevance of a passage may be marked on a binary (e.g. Relevant or Not Relevant) or graded (e.g. Perfectly Relevant, Partially Relevant, Related, Irrelevant) scale. In CURE passages are marked as Relevant, Partially Relevant and Not Relevant with respect to each query.
@@
CURE includes $20002000$ queries evenly distributed over 10 splits corresponding to the following domains: Dentistry and Oral Health, Dermatology, Gastroenterology, Genetics, Neuroscience & Neurology, Orthopedic Surgery, Otorhinolaryngology, Plastic Surgery, Psychiatry & Psychology and Pulmonology. Each split contains $5050$ layman and $150150$ expert level queries.
@@
In addition to identified relevant passages, CURE’s corpus includes passages mined from an index of over 7777M biomedical passages. These were extracted from articles whose publishers permit their use for evaluation. Sources include the PMC Open Access Subset of Medicine as well as articles scraped from Nature and BioMedCentral. Articles were parsed using the JusText parser and segmented into passages using the parsed document structure. Low-quality passages were filtered out by employing a selection of heuristics introduced to clean training data for Gopher Rae et al. (2021). The corpus contains 244600 passages from 51083 articles. The distribution of passages and articles across domains is presented in Table 5.
@@
CURE was built in three phases: Annotation, Query Translation and Corpus Construction.
@@
During the Annotation phase, medical practitioners expressed their information needs as queries and identified relevant passages in literature the consulted. Annotations were then validated and processed.
@@
During query translation, queries generated by annotators were translated into English(en), French (fr) and Spanish (es) as needed.
@@
After Corpus Construction, the English queries were used to construct the monolingual conditions of CURE while the Spanish and French queries were used to construct the cross-lingual conditions.
2.1 Annotation
To ensure high quality annotations, CURE was constructed with medical professionals. The annotation team included doctors and nurses that were recruited from our professional networks, as well as native speakers of French, English and Spanish. Annotators were compensated for their work. To ensure consistency and quality, annotations were validated and processed. Annotation was carried out over a 9 month period by twelve annotators.
@@ 2.1 Annotation
To create a test dataset with queries representative of the information needs of point-of-care medical practitioners, we asked annotators to formulate queries representing their own information needs and identify relevant passages from biomedical articles they consulted to meet these needs. Annotators were instructed to express queries as natural language questions in English, French or Spanish choosing the language they felt most comfortable with. Annotators identified journal and articles they felt were appropriate given their needs and specialties and we ensured that publishers that permitted their use in evaluation prior to annotation. A copy of our annotation instructions is presented in section 9. An example of a generated question and identified relevant passage is presented in Table 1.
@@
Table 1: Example of a generated query and identified relevant passage from the domain of Plastic Surgery
Argilla was used to collect annotations. The interface for annotation is presented in section 9. Annotators provided the text of the query and of identified relevant passages. Additionally, they were asked to specify the query’s language, discipline/s and its level of expertise. Annotators identified queries as being layman or expert with the former being representative of those posed by a medical practitioner seeking to refresh their knowledge or communicate with laymen, and the latter being representative of queries posed by a specialized medical practitioner. An example of a layman and expert query is presented in Table 2.
@@
Table 2: Example of layman query and expert query in the domain of Dental and Oral Health
@@
Lead annotators from the annotation team validated annotations, ensuring they were of a high quality. In evaluating generated queries, they considered the following criteria:
@@
  • Does the question match the level of expertise it is associated with?
  • Is the question well-written? Considering its syntax, fluency and presence of typos.
@@
Identified relevant passages were also validated to ensure that they were self-contained and met the information need expressed by the query.
If annotations failed to meet requisite criteria, validators provided feedback and corrected annotations were reevaluated prior to inclusion in the dataset. An example of an invalid annotation with corresponding feedback is presented in Table 3.
@@
Table 3: Example of annotation with feedback provided during validation
@@
Annotators may have entered slight conditions of a single passage as relevant for distinct queries as a passage may meet more more than one information need and the text of relevant passages were entered manually. Furthermore a passage that was identified as relevant for one query may have been embedded in a passage identified as relevant for another query. Although not identified as such, the containing passage would be relevant with respect to the query associated with the contained passage.
In processing annotations, minor variants of the a single passage were identified as such and a representative passage was selected to replace variants. Additionally containing passages were identified and marked as partially relevant for the query associated with the contained passage.
2.2 Query Translation
Using DeepL API, queries were translated so each query is available in English, French and Spanish. DeepL was selected based on its performance relative to alternative tools on a small test of translations evaluated by native speakers. Evaluation criteria included: accuracy, clarity, fluency, appropriateness, consistency, formatting, preservation of proper names and terms, grammar and style, and formal correspondence. After, native speakers performed spot checks to ensure translations were of quality.
2.3 Corpus Construction
Our corpus was constructed by mining for in domain passages and assigning relevance judgments to passages from large candidate pools using an LLM.
@@ 2.3 Corpus Construction
Following Zhang et al. (2023), we mine passages using Reciprocal Rank Fusion (RRF) combining KNN and BM25 retrievers. Iterating through all validated queries, a search was executed using the query and one relevant passage with each type of retriever. Embeddings for the KNN retrievers were generated using BGE-small-v1.5 Xiao et al. (2023). Each KNN retriever considered 250 passages from each shard and returned 50 passages. The BM25 retriever using the query employed a match query with a slop of 2 while the BM25 retriever using the relevant passage employed a match phrase query permitting fuzzy matches. Source articles contributing to the top 50 passages returned were considered to be in-domain and their passages were included in CURE’s corpus.
@@
In any sizable retrieval dataset, it is not possible to manually annotate all query-passage pairs for relevance. Typically, small sets of candidates are selected for annotation while the remaining passages are presumed to be irrelevant. As Thakur et al noted, in constructing BEIR, since only small candidate pools are annotated for relevance, datasets can have a bias. For example, Thakur et al. identified a bias in TREC-COVID likely stemming from a use of lexical cues to construct candidate pools Thakur et al. (2021).
To mitigate this bias, we generated large candidate pools for each query using the GTE multi-lingual cross-encoder Zhang et al. (2024). Generated queries and associated passages were assigned to distinct splits, with each split corresponding to one of 10 medical domains. Queries and passages belonging to more than one domain were placed in only one split. Each english query was scored against all passages within the split, and the pool for each query was constructed by taking passages with a score above 0.42. This cut-off score was established using a small test set of passages annotated by humans for relevance and was found to maximize the number of relevant passages included in the pool while minimizing the number of irrelevant passages.
@@
Passages in candidate pools were annotated for relevance using Qwen 2.5 72727272B Yang et al. (2024). The use of LLMs to annotate these pools of passages was inspired by Upadhyay et al.’s work, which found a strong correlation between human relevance judgments and LLM derived judgments Upadhyay et al. (2024). Qwen 2.5 was selected due to its relatively strong concordance with a human on small test set. We used a variation of the prompt used by Upadhyay et al. Generated explanations are included in the dataset. An example explanation is presented in Table 4.
@@
Table 4: Example of a relevance judgment with explanation
After the construction of the corpus, query translations were used to generate a monolingual (English) and two cross-lingual conditions (Spanish-English, French-English) of the dataset with conditions differing only in query language.
3 Dataset Statistics
The average number of tokens per query and passages in the corpus are presented in table Table 5.

### v4

@@ Abstract.
Given the dominance of dense retrievers which do not generalize well beyond their training dataset distributions, domain-specific test sets are essential in evaluating retrieval performance. Few test datasets are available for retrieval systems intended for use by healthcare providers in a point-of-care setting. To fill this gap we have collaborated with medical professionals to create CURE, a test dataset for passage retrieval composed of 2,000 expert written queries spanning 10 medical domains with a monolingual (English) and two cross-lingual (French/Spanish to English) conditions. In this paper, we describe how CURE was constructed and provide baseline results to showcase its effectiveness as an evaluation tool. CURE is published with a CC BY-NC 4.0 license and can be accessed on Hugging Face [^1] https://huggingface.co/datasets/clinia/CUREv1 and as a retrieval task on MTEB.
Biomedical information retrieval, Cross-lingual retrieval, Test datasets, Point-of-care, Dense retrievers, Sparse retrievers
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737435isbn: 979-8-4007-1454-2/2025/08ccs: Information systems Information retrievalccs: Information systems Environment-specific retrievalccs: Information systems Test collectionsccs: Information systems Retrieval effectivenessccs: Information systems Presentation of retrieval results
1. Introduction
@@
Evidence-based medicine requires point-of-care healthcare professionals to stay up to date with advances in their specialties to make data-driven decisions when providing care to patients (Sackett, 1997). Despite this imperative, research findings have been estimated to take up to 17 years to fully implement in practice (Balas and Boren, 2000).
@@
An impediment for practitioners is the increasingly rapid rate at which relevant literature is published (Kamtchum-Tatuene and Zafack, 2021). Each year, more than one million new articles are published in biomedicine and life sciences (González-Márquez et al., 2024).
@@
Given the ever-expanding landscape of biomedical information, applications that support ad hoc retrieval are essential tools for health care professionals to provide evidence-based care (Lu, 2011). Using ad hoc retrieval, a healthcare professional can express their information need as a query, initiating a search that returns likely relevant documents from a vast corpus (Lu, 2011).
@@
Although the vast majority of published biomedical literature is in English, healthcare providers may practice in other languages (Baethge, 2008). To effectively accommodate this linguistic diversity, tools aimed at efficient biomedical information access should enable cross-lingual retrieval.
@@
Information retrieval systems and techniques are typically evaluated using test datasets. There are a wide variety of general test datasets available for evaluating retrieval. At its inception, MTEB (Massive Text Embedding Benchmark) made available 15 retrieval test sets, a collection that continues to expand (Muennighoff et al., 2022). To our knowledge, there are no cross-lingual datasets suitable for evaluating retrieval systems intended for use by healthcare providers in a point-of-care setting. As such, domain-specific test sets are essential given the dominance of dense retrievers, whose performance does not generalize well to out-of-domain tasks (Thakur et al., 2021; Ren et al., 2023).
@@
At Clinia, we build products that power search and discovery in health workflows, and the absence of such test datasets has hampered product development, as we cannot benchmark developed solutions systematically. To address this gap, we have introduced a domain-specific test set, which we have released under an open-source license.
@@
In this work, we introduce Clinical Understanding & Retrieval Evaluation (CURE), a test dataset to evaluate ad hoc retrieval systems for use by medical practitioners in a point-of-care setting. We present results from competitive baselines to demonstrate the utility of CURE as a test dataset and to enable comparison with future systems.
@@
CURE is a retrieval dataset with a monolingual and two cross-lingual conditions, with splits spanning ten medical domains. Queries in CURE are natural language questions formulated by health care providers. They express the information needs of practitioners consulting academic literature in the course of their duties. Queries are available in English, French and Spanish. The corpus is constructed by mining an index of English passages extracted from biomedical academic articles. We have made this dataset available to spur research and development on retrieval in a point-of-care setting.
2. Dataset Construction
@@
CURE is a test collection for passage retrieval. In passage retrieval, given a corpus C of n passages and a query q, the objective is to retrieve a list of k passages, sorted by decreasing relevance. A relevant passage meets the information need expressed by the query. In the context of a dataset, the relevance of a passage may be marked on a binary (e.g. Relevant or Not Relevant) or graded (e.g. Perfectly Relevant, Partially Relevant, Related, Irrelevant) scale. In CURE, passages are marked as Relevant, Partially Relevant, and Not Relevant for each query.
@@
CURE includes 2,000 queries evenly distributed over 10 splits corresponding to the following domains: Dentistry and Oral Health, Dermatology, Gastroenterology, Genetics, Neuroscience & Neurology, Orthopedic Surgery, Otorhinolaryngology, Plastic Surgery, Psychiatry & Psychology, and Pulmonology. Each split contains $5050$ layman and $150150$ expert level queries.
@@
In addition to identified relevant passages, CURE’s corpus includes passages mined from an index of more than 7777M biomedical passages. These were extracted from articles whose publishers allowed their use for evaluation. Sources include the PMC Open Access Subset (of Medicine, 2003) as well as articles scraped from Nature and BioMedCentral. Articles were parsed using the JusText parser and segmented into passages using the parsed document structure. Low-quality passages were filtered out by employing a selection of heuristics introduced to clean training data for Gopher (Rae et al., 2021). The corpus contains 244,600 passages from 51,083 articles. The distribution of passages and articles across domains is presented in Table 6.
@@
CURE was built in three phases: Annotation, Query Translation, and Corpus Construction.
@@
  • During the annotation phase, medical practitioners expressed their information needs as queries and identified relevant passages in the literature they selected. The annotations were then validated and processed.
@@
  • During query translation, queries were translated from the language in which the annotators had originally written them into English (en), French (fr) and Spanish (es) as needed.
@@
After corpus construction, English queries were used to construct the monolingual conditions of CURE, while Spanish and French queries were used to construct cross-lingual conditions.
2.1. Annotation
To ensure high-quality annotations, CURE was constructed with medical professionals. The annotation team included physicians and nurses who were recruited from our professional networks, as well as native speakers of French, English, and Spanish. Annotators were compensated for their work. To ensure consistency and quality, the annotations were validated and processed. Annotation was carried out over 9 months by 12 annotators.
@@ 2.1. Annotation
To create a test dataset with queries representative of the information needs of point-of-care medical practitioners, we asked annotators to formulate queries based on their own information needs and identify relevant passages from biomedical articles they selected to address those needs. Annotators were instructed to express their queries as natural language questions in English, French, or Spanish, choosing the language they felt most comfortable with. Annotators identified journals and articles they deemed appropriate for their specialties and needs. Before annotation, we ensured that publishers permitted the use of these articles for evaluation.
Annotators were provided with a detailed set of guidelines with accompanying examples to help them with their task. Below is a condensed version for reference:
Using a selected scientific article of your choice, generate a query-passage pair and evaluate its audience.
  • The passage should either be or contain the answer to the query and should be an excerpt from the article. Do not paraphrase.
  • The passage should be a single paragraph or a short list.
  • If a single query is answered by multiple passages, please generate distinct query-pairs for each passage.
  • Identify the article by providing its URL and its title.
  • If available, provide the full text of the article.
An example of a generated question with the corresponding identified relevant passage is presented in Table 1.
@@
Table 1. Example of a generated query and identified relevant passage from the domain of Plastic Surgery
Argilla (Vila-Suero and Aranda, 2023) was used to collect annotations. The annotation interface is presented in Appendix A. Annotators provided the query text and identified relevant passages. Additionally, they specified the query’s language, discipline(s), and level of expertise. Annotators identified queries as layman or expert. Layman queries represent those asked by medical practitioners seeking to refresh their knowledge or communicate with non-experts, while expert queries reflect those posed by specialized medical practitioners. An example of a layman and expert query is presented in Table 2.
@@
Table 2. Example of layman query and expert query in the domain of Dental and Oral Health
@@
Each annotation was reviewed by lead annotators to ensure high quality. In evaluating generated queries, they considered the following criteria:
@@
  • Does the question match the level of expertise with which it is associated?
  • Is the question well written, considering syntax, fluency, and the presence of typos?
@@
The identified relevant passages were also validated to ensure that they were self-contained and met the information need expressed by the query.
If annotations did not meet the required criteria, the validators provided feedback and corrected annotations were re-evaluated before inclusion in the dataset. Across all annotations, 12.4% of annotations were flagged for revision. The most common sources of error were related to formulation. An example of an invalid annotation with the corresponding feedback is presented in Table 3.
@@
Table 3. Example of annotation with feedback provided during validation
@@
Annotators may have entered slight variations of the same passage as relevant for distinct queries as a passage may meet more than one information need. Additionally, since relevant passages were entered manually, variations in text may have occurred. Furthermore, a passage that was identified as relevant for one query may have been embedded in a passage identified as relevant for another query. Although not identified as such, the containing passage would still be relevant for the query associated with the contained passage.
In processing annotations, minor variants of the same passage were identified automatically using Levenshtein distance between passages as a heuristic and the longest passage among variants was selected to replace variants. In addition, passages containing these variations were identified and marked as partially relevant for the query associated with the contained passage.
2.2. Query Translation
Queries were translated so that each query is available in English, French, and Spanish. To select the best-performing translation tool, we sampled a set of queries and ran them through various candidate tools. Then, three native-speaking medical annotators were asked to rate each translation from 0 (not meeting criteria) to 2 (fully meeting criteria) based on how well they met each of the following criteria:
  • Accuracy: The content should faithfully convey the message and sentiment of the original text in the target language. No added or forgotten content.
  • Clarity: The message must be easily understandable, and any instructions should be actionable and easy to follow.
  • Fluency: Does the translation flow naturally in the target language?
  • Appropriateness: Consider cultural nuances and ensure that the translation is appropriate for the target audience. Check for register, idiomatic expressions and cultural references that may need adaptation.
  • Proper names and terms: Are medical terms, names, trademarks, and other non-translatable words preserved from the source text?
  • Grammar and style: Check for grammatical correctness in the translated text. Evaluate whether the translator has captured the style and tone of the original content.
Averaged results across annotators and queries for each tool are presented in Table 4
Model Name Accuracy Clarity Fluency Appropriateness Proper Names Grammar
Mistral-7b-Instruct (Jiang et al., 2023) 1.1 1.133 1.2 1.833 0.967 1.3
Bing Microsoft (Microsoft, 2024) 1.767 1.7 1.567 1.933 1.733 1.733
ChatGPT4-turbo-2024-04-09 (OpenAI, 2024a) 1.933 1.933 1.8 1.9 1.767 1.967
Reverso (Reverso, 2024) 1.833 1.733 1.767 1.833 1.6 1.867
Facebook-m4t-v2-large (Seamless Communication, 2023) 1.6 1.567 1.667 1.933 1.467 1.833
Llama-3-8b-instruct (AI@Meta, 2024) 1.6 1.433 1.367 1.933 1.233 1.467
DeepL (SE, 2024) 1.933 1.933 1.833 1.967 1.867 1.967
Table 4. Evaluation Results for Translation Tools
In selecting a tool, the criteria accuracy and proper names and terms were given the highest weights, followed by clarity and fluency. The criteria of appropriateness and grammar and style were given the lowest weights. DeepL API (SE, 2024) was selected based on its strong performance relative to alternative tools, and we used it to translate the remaining queries. Finally, native speakers performed quality checks and made manual modifications to ensure that the resulting translations were of high quality.
2.3. Corpus Construction
Our corpus was constructed by mining for in-domain passages and assigning relevance judgments to passages from large candidate pools using an LLM.
@@ 2.3. Corpus Construction
Following Zhang et al. (2023), we mined passages using Reciprocal Rank Fusion (RRF) combining KNN and BM25 retrievers. Iterating through all validated queries, a search was executed using each query and one relevant passage with each type of retriever. Embeddings for the KNN retrievers were generated using BGE-small-v1.5 (Xiao et al., 2023). Each KNN retriever considered 250 passages from each shard and returned 50 passages. The BM25 retriever using the query to search employed a match query with a slop of 2 while the BM25 retriever using the relevant passage to search employed a match phrase query permitting fuzzy matches. Source articles contributing to the top 50 passages returned were considered to be in-domain and their passages were included in CURE’s corpus.
@@
In any large retrieval dataset, manually annotating all query-passage pairs for relevance is not feasible. Typically, small sets of candidates are selected for annotation while the remaining passages are presumed to be irrelevant. As Thakur et al. noted, in constructing BEIR, since only small candidate pools are annotated for relevance, datasets can be biased. For example, Thakur et al. identified a bias in TREC-COVID likely arising from the use of lexical cues to construct candidate pools (Thakur et al., 2021).
To mitigate this bias, we generated large candidate pools for each query using the GTE multilingual cross-encoder (Zhang et al., 2024). The generated queries and associated passages were assigned to distinct splits, each corresponding to one of ten medical domains. Queries and passages belonging to multiple domains were placed in a single split. The English variant of each query was scored against all passages within the split, and the pool for each query was constructed by taking passages with a score above 0.42. This cutoff score was established using a test set of passages annotated by humans for relevance and was found to maximize the difference between the rate of relevant passages and non-relevant passages included in the candidate pool.
@@
The passages in the candidate pools were annotated for relevance using Qwen 2.5 72727272B (Yang et al., 2024). The use of LLMs for passage annotation was inspired by Upadhyay et al.’s work, which found a strong correlation between human relevance judgments and LLM-derived judgments (Upadhyay et al., 2024). Qwen 2.5 was selected because of its relatively strong concordance with human annotations on a test set. We used a variation of the prompt from Upadhyay et al presented in Appendix B. Generated explanations are included in the dataset. An example explanation is presented in Table 5
@@
Table 5. Example of a relevance judgment with explanation
. To evaluate the quality of relevance judgments in our final dataset, we sampled a set of query-passage pairs from our dataset and had two annotators independently assess relevance. Their judgments were compared with those in our dataset, yielding a Fleiss’ Kappa of 0.691.
After constructing the corpus, query translations were used to generate a monolingual (English) and two cross-lingual conditions (Spanish-English, French-English) of the dataset with conditions differing only in query language.
3. Dataset Statistics
The average number of tokens per query and passages in the corpus is presented in table Table 6.